South Pole Diaries 1997/98

   

   


Thursday 27th November 1997 - Disaster!

From John Storey.....

Today probably would have been a good a day if we hadn't tried to take some data with the Mism. It quickly became clear that there was a large signal no matter where in the sky we looked, and that although the signal varied with postion it was not in the way it was supposed to. A little investigation, with Ant waving first his hand and then a soldering iron around in front of the window, showed that only one beam was present. That gave two possibilities, both ugly: either something was blocking one of the beams, or one of the mirrors had fallen off. Given the glass-rattling-about noises we heard the other night when adjusting the rotator, the latter and uglier possibility seemed the more likely.

We removed the optics plate - a nasty operation to perform in the cold because the plate has to be lowered an inch or two and then held while the D-connector is unplugged and the window heater wires unscrewed from the terminal block. We put the optics plate in a plastic bag and allowed it to warm up inside the AASTO for a few hours.

It turns out that one of the little rectangular beam-steering mirrors has fallen off. Of the three dabs of epoxy holding it in place, one has simply dropped off both the glass and the aluminium (it appears not to haave properly adhered in the first placce), one has remained attached to the aluminium plate and the glass but has pulled a big (5mm) chip of glass out of the back of the mirror, and the other has come off the aluminium *and* also pulled a chip of glass off the mirror.

The possibilities now seem to be:

1. Bring the optical boxes back to Sydney and make proper optical mounts, returning the boxes in January. This is probably not practical as we would need complete new sets of mirrors. It's something we certainly should do before the end of next year, though.

2. Re-glue the mirror, using cryogenically rated epoxy. If we take the second option, we need to realign the optical system with a laser. This will be time-consuming, but possible. What we would do is position a piece of paper above the optics plate at the same height as the outer surface of the window. Then, set up a laser so that its beam passed through the "intact" channel and comes out through the centre of the window. Finally, align the repaired mirror so that its beam crosses the first one at the paper. Sound ok?

Some questions for Max:

1. What is the f number and diameter of the beam as it passes through the window? 2. What epoxy was used the first time? 3. How critical is the alignment (we've found all the little shims - we think)

We'd like to think about it for a day before we proceed - advice welcome!

Some other software issues came up with the Mism:

1. "cvfma" throws the cvf into continuous rotation. Is cvfma in steps or degrees? Only "steps" is meaningful, of course. 2. The figures given to convert rotator steps into degrees appear to be correct. A quick "eyeball" calibration gives:

ccw beam on horizon ~5,000 steps cw beam at zenith ~15,000 steps box horizontal ~23,000 ccw beam at zenith 30,000 cw beam at BB 37,000 (assuming of course that there is a ccw beam, which there isn't)

Can I stop being depressed now?

It's possible to cheer yourself up a bit by looking at:

http://141.224.128.11/ago/daily_reports/ago_today.lis

where you'll see the AASTO data from the ARGOS transmissions (listed as AGO-A2). All the numbers are complete nonesense until we send Augsburg the correct calibration file. (Actually, the wind speed and air pressure are correct, and the TEG volages are nearly right).

With mcba's help we got the 386 super set up so that people not at the South pole can talk to it. It turns out that the gateway and domain addresses on the hard disc had not been configured for this location. The 386 super still crashes (or hangs) if a telnet command has an unsatisfactory result - eg if it tries to talk to an instrument that's not there.

Before firing up the Nism again we tracked down the faulty Dallas problem and checked the limit switches again. I was worried that whatever had killed the Dallas (assuming it wasn't Larry Hagman) might be the same thing that caused the limit switch to go non-functional a few days back. Anyway, the limit switches are fine (heaps of megohms, says the Fluke) and the Nism fired up without any problems.

Here's a rough calibration of where things are in absolute degrees (ie, using "rotma")

ccw beam peaks at BB ~175 deg (abs) box horizontal ~390 cw beam hits snow ~670

There's a good strong signal that does all the right things. However, looking at the adc readings it's very noisy. Even when the signal has a s/n ratio of 100:1, the output from "det" and "detx" is all over the place. I hope the data acquiistion software is making multiple samples.

"adc amps" also gives a very noisy result, although this cleans up when the Stirling cooler is off. It's likely that the current *is* in fact noisy.

"adc volts, bb, cooler all give very stable output with fluctuations of between 2 and 5 ADU.

The faulty Dallas turned out to be just that. It was the one on the Nism mounting bracket. Even after I cleaned it in bicarb soda, washed it, dried it and brushed it, there's still no response. It's been replaced with another one, which the software correctly identifies as "no name found". (The Dallas whose name can not be spoken.)

Occasionally we get "sbit" as a response when we do a DS INIT, but we're not losing any sleep over it.

When I was wrestling with the Dallas I kicked the bucket of sodium bicarbonate over, and now it's all over the AASTO. (But it's a hell of an improvement on either brown slime or rockwool, believe me.) There's an outside chance that when we come back to the AASTO this time next year it will have grown stalagtites and stalagmites, and be really quite attractive.

The evening picked up considerably when we helped Fred unbolt the 300lb friction disc from SPIREX. Yes, Fred had lifted the crane into position single-handedly last night, while we were asleep. Once we got the disk down, Fred lifted it onto the trolley and wheeled it across the roof of MAPO. (It's not clear now why we went to the trouble of installing a crane. It would have been much simpler just to bolt Fred to the platform.)

Preparations for Thanksgiving are now in full swing (It's celebrated on Saturday at the South Pole - I'm not sure what time zone that corresponds to!) The Dome was full of smoke this evening and no-one was worried - the cooks had just built a turkey smoker and fired it up (actually, given the beer situation, I suspect that home-brew bourbon might also be part of the plan).

Abu is still on the pump; the pressure is slowly going down, as well it should. Ant leak-checked the T-tube and found it leak-tight, which is a worry because everyone knows that fluorosilicone O-rings are slightly permeable to helium. Ant is going to try a different kind of leak checker, and possibly a different kind of helium. Actually, Ant just came in very pleased and said yes, he *had* got the O-rings to diffuse helium. Good work, Ant. He's also replaced the "warm" O-rings on the T-tube with Viton ones. They be fine as long as the heaters don't fail - in which case Abu will leak anyway because it contains most of the entire Parker O-ring inventory.

I visited the ham shack in pursuit of someone who could make up a 25-foot length of 10-base T cable for me, and found the operator in the midst of a series of 30-second conversations with a bunch of anonymous people. He seemed to be enjoying himself, though I'm not sure why. If I get a chance I'll fire the rig up and see what's out there.

Nothing else good happened today.

John

 

 

Information

Further Information

Contact: